A sinusoidally varying voltage is represented by

0 0
Answer #1

Comparing given equation with standard equation:

V(t) = V_max*sin (w*t)

V_max = Peak Voltage

w = Angular frequency = 2*pi*f

f = frequency = w/(2*pi)

Now From above given equation:

V(t) = (62.5 V)*sin (17071t)

So

Part A

frequency = f = 170*pi/(2*pi) = 85 Hz

Part B.

V_max = Peak Voltage = 62.5 V

Please Upvote.